# An Empirical Study of Cross-Lingual Transfer Learning in Programming Languages

### v2

@@ IV-B Performance Prediction
TABLE I: Features of language pairs. Column “Measure” indicates how the feature is computed: (e)quivalence, (o)verlap, (s)ource, (t)arget, (b)oth source and target, (rd) for relative difference. For dataset specific features, ($x$) can be s, t, or rd. Details of the linguistic features are available in supplementary material Tab. IX
Given a task and target language, how should we pick the source language for best performance? Knowing this without having to train many models on different source languages can save significant resources. It also provides a basis to study important features that characterize a successful transfer from a language to another. To predict the performance of the models on the different tasks for different source languages, we train a ranking model based on gradient-boosted decision trees (GBDT) [27]. We show that our models outperform baselines based on regressors or programming-language history. We consider the features defined in Tab. I. These features can be grouped into four categories.
  1. 1.
    Linguistic features: general properties characterizing a language pair like whether they support the same paradigms, the same style of memory management (is there a garbage collector or not), or the same type-system characteristics.
  2. 2.
    Syntactic features: properties of the syntax of a language pair measured by overlap in the counts of certain token types, as determined using the Pygments library on the corpus of code in the given languages.
  3. 3.
    Dataset-specific features: properties of the problems associated with code samples for the language pair in the online coding judge website from which the dataset came.
  4. 4.
    Model-specific features: whether the model saw the source language, target language, or both during pretraining.
@@
To predict the top source languages for a given task and target language $L_{t}$ in the set of target languages $T$, we train a ranker using the LightGBM [27] implementation of the LambdaRank algorithm [28]. The model takes features in Tab. I as inputs and scores source languages $L_{s}$ in terms of their relevance to the target $L_{t}$. Our ranker uses a boosting ensemble of 100 decision trees with 16 leaves each. We consider the normalized discounted cumulative gain score (NDCG@K) at $K=3$ as our evaluation metric. We evaluate the model using leave-one-out (LOO) cross-validation on the set $T$. For each target language $L_{t}$, we train a ranker to predict rankings of different sources for each language in $T$ leaving out the source ranking for $L_{t}$ as a test set. For each fold, we compute the NDCG@3 score on the test set. We compare the performance of our ranker with a regression-based ranker and a history ranker. The history ranker selects the top $K$ source languages with the shortest path to the given target language in a programming languages history graph [^2] http://rigaux.org/language-study/diagram.html (contracted to merge different versions of the same language). The regression-based ranker uses a LightGBM regressor with the same hyperparameter settings as our ranker and selects the top $K$ source languages with the highest predicted scores. Fig. 6 shows the mean and standard deviation of NDCG@3 using LOO on $T$. While the error bars are overlapping, both regression and rankers outperform the history ranker on all tasks, indicating the importance of considering a variety of features for predicting top-performing sources rather than simplified heuristics. Our model outperforms baseline rankers on all tasks with the exception of clone detection, for which it shows slightly lower performance than the regression model. This indicates that ranking is the best suited method for predicting top performing sources. Based on these results, we use our ranker for the feature analysis in RQ3.
@@
Figure 6: NDCG@3 scores for different rankers and tasks corresponding to LOO evaluation over the set of target languages.
Using a ranking model and considering a variety of features rather than simplified heuristics is crucial for predicting top performing sources.
IV-C Feature analysis
@@
Figure 7: Normalized SHAP values aggregated by tasks for the features defined in Tab. I. The features are sorted by mean rank. Grayed-out features were unavailable for the given task.
@@
While Sec. IV-A explored how well transfer learning performs for different languages on different tasks, we would like to dig deeper into why they perform that way. That is, how do the features of language pairs affect transfer? Which characteristics of the source and target language best explain the variation of transfer across tasks, the dependence of transfer on source and target pairing, as well as the superior transfer supported by specific sources? These questions can be answered by measuring feature importance.
@@
Unfortunately, most traditional techniques for measuring feature importance suffer from a common pitfall: when two features overlap, they arbitrarily assign more importance to one at the expense of the other. The state-of-the-art technique for measuring feature importance avoids this pitfall by considering all possible subsets $S\subseteq F$ of features [29]. Then, the importance of a given feature is the difference it makes to model performance in combination with all possible sets of other features. The resulting SHAP values [29] show the relative impact of each feature on the model output. In the context of ranking different source languages for cross-lingual transfer, SHAP values measure how different features contribute to transfer. Fig. 7 shows SHAP values computed based on the ranking models evaluated in Sec. IV-B. For presentation purposes, it rescales the values so the most important feature for each task has importance 1.
Task-dependent feature importance
@@
While previous works emphasize the importance of specific features for cross-lingual transfer, the lack of comparison on different tasks limits their conclusions. The supported cross-task analysis demonstrates a task-dependent feature importance whereby different features contribute differently to cross-lingual transfer depending on the task. For example, the top features for the solution domain classification task are Difficulty, Literals, Names, and Operators. Being able to predict the solution domain for a code solution requires some knowledge of the underlying problem for which the tags are an attribute. The difficulty score is another attribute of the problem. Problems with similar difficulty scores require similar algorithms. The literals, names, and operators are other indicators of the algorithms used in a code sample. In comparison, the most significant features for the clone detection task are Token, Names, and Keywords. Detecting a clone requires different skills than classifying domains. A deeper understanding of the code semantics irrespective of the similar algorithms used is needed. The overlaps in names, keywords, and more generally tokens is key for understanding the semantics of code [11].
Range of important features
@@
Different tasks seem to not only focus on different features, but also focus on a different number of features. While the feature importance for solution domain classification and error detection is concentrated on a small number of features, the feature importance is more spread out for clone detection and code repair. For example, the code repair task requires a transfer of knowledge from overlapping keywords, names, and more generally different tokens from a source language for fixing a code in a transfer language. On the other hand, solution domain classification focuses on fewer features, related to the problems attributes.
Comparison across categories
@@
Comparing the importance of different feature categories across tasks, we observe a relatively higher significance of syntactic features as demonstrated by the top two features Keywords and Names. Compared to transfer learning across natural languages [17], model-specific and linguistics features seem to carry less significance for transfer in programming languages. While de Vries et al. [17] find the potential of cross-lingual transfer dropping significantly for sources and targets that are not seen by the model, our finding brings hope to not only transferring learning to low-resource target languages, but even finetuning the model on unseen languages that could serve as good sources, such as Kotlin.
Within-category differences
@@
While different categories seem to have a higher impact on the model predictions, the impact of single features within a category varies across task. Our results are consistent with prior work that found Names and Keywords to be the most important features for cross-lingual transfer in certain tasks [11]. However, half of the tasks show keywords to be more important than names, while the other half has the opposite order of importance.
1. The importance of language pair features for cross-lingual transfer performance depends on the task. 2. Higher-level tasks seem to draw upon a broader range of features to best understand the code. 3. Overall, the most important features across the four tasks studied in this paper are keywords and names.
V Threats to Validity
V-A External validity
One threat is that to make the large number of experiments feasible and reduce their carbon footprint, we had to limit ourselves to a moderately-sized model. We mitigate this threat by picking CodeT5 [22], which is widely used in the research literature for AI-for-Code. We believe that even if we had used larger models, insights from one model would not perfectly generalize to another.
Another threat is that to cover many languages, we restricted ourselves to four tasks. We mitigate this threat by picking tasks that cover a broad spectrum of difficulty and are diverse, as evidenced by their different performance and feature importances. Overall, our number of tasks exceeds prior work on cross-lingual transfer for programming languages.
@@ V-A External validity
A further threat is that our datasets have some synthetic aspects, such as type-IV clones [25] for clone detection data or fault injection for code repair data. However, using fully natural data is such a high standard that hardly any academic paper satisfies it, nor is it typically the case in industry practice. Instead, we mitigate this threat by explicitly measuring the importance of dataset-specific features of language pairs.
@@
Another threat is that we focused on the setting of zero data in the target language. This represents the lowest barrier to entry, as gathering even a small number of examples can be tricky. Therefore, we leave exploring the case where there is some target-language data to future work.
V-B Internal validity
@@
We believe that our study avoids one common internal validity threat, which revolves around insufficient runs. Thanks to the sheer number of language pairs for each task, we get thousands of different experiments, whose results are consistent with each other. Moreover, each of these experiments involves fine-tuning with tens of thousands of samples, giving us confidence that the choice of samples is not a threat either.
Since our study is the first of its kind at this scale, there is a threat that it cannot be directly compared to baselines. We mitigate this by observing that our results do confirm more narrow studies where applicable, for instance, a study of transfer for identifier-centric tasks [11].
VI Conclusion
We perform a systematic and extensive study of LLM transfer learning covering up to 41 programming languages, across 4 tasks including both classification and generation. To understand relative differences in cross-lingual performance between different programming languages, we define several linguistic, syntactic, dataset, and model-specific features of language pairs and then analyze the feature importance for a model that predicts transfer performance. Overall, we hope that this paper helps advance the community’s understanding of how learning transfers among programming languages, and that this improved understanding leads to better models to assist users of those languages, particularly low-resource ones.
VII Data Availability
The experiments are based on the publicly available CodeT5-base (220M parameters) model [22] and the open-sourced datasets CodeNet [19] and XCodeEval [20]. Upon paper acceptance, we plan to open-source the code to create the derived datasets and the framework to train and evaluate the models. Also, we plan to make the 58 finetuned models available with the final version of the paper.
We also plan to submit a replication package to the artifact committee.

### v3

@@ 4.2 Performance Prediction
Table 9: Features of language pairs. Column “Measure” indicates how the feature is computed: (e)quivalence, (o)verlap, (s)ource, (t)arget, (b)oth source and target, (rd) for relative difference. For dataset specific features, ($x$) can be s, t, or rd. Details of the linguistic features are in Table LABEL:tab:linguistic-features.
@@
To predict the top source languages for a given task and target language $L_{t}$ in the set of target languages $T$, we train a ranker using the LightGBM (Ke et al., 2017) implementation of the LambdaRank algorithm (Burges, 2010). The model takes features in Table 9 as inputs and scores source languages $L_{s}$ in terms of their relevance to the target $L_{t}$. Our ranker uses a boosting ensemble of 100 decision trees with 16 leaves each. We consider the normalized discounted cumulative gain score (NDCG@K) at $K=3$ as our evaluation metric. We evaluate the model using leave-one-out (LOO) cross-validation on the set $T$. For each target language $L_{t}$, we train a ranker to predict rankings of different sources for each language in $T$ leaving out the source ranking for $L_{t}$ as a test set. For each fold, we compute the NDCG@3 score on the test set.
We compare the performance of our ranker with a regression-based ranker and a history ranker. The regression-based ranker uses a LightGBM regressor with the same hyperparameter settings as our ranker and selects the top $K$ source languages with the highest predicted scores. The history ranker selects the top $K$ source languages with the shortest path to the given target language in a programming languages history graph [^2] http://rigaux.org/language-study/diagram.html (contracted to merge different versions of the same language). Figure 7 shows the mean and standard deviation of NDCG@3 using LOO on $T$. While the error bars are overlapping, both regression and rankers outperform the history ranker on all tasks, indicating the importance of considering a variety of features for predicting top-performing sources rather than simplified heuristics. Our model outperforms baseline rankers on all tasks with the exception of clone detection, for which it shows slightly lower performance than the regression model. This indicates that ranking is the best suited method for predicting top performing sources. Based on these results, we use our ranker for the feature analysis in RQ3.
@@
Figure 7: NDCG@3 scores for different rankers and tasks corresponding to LOO evaluation over the set of target languages. The first four tasks are evaluated on ranking predictions on CodeT5 transfer experiments, while code generation is evaluated over ranking predictions on Llama 3.3 70B-Instruct.
Rankers outperform regressors for predicting top-performing sources. Using historical relationships among programming languages alone is insufficient.
4.3 Feature analysis
@@
Figure 8: Normalized SHAP values aggregated by tasks for the features defined in Table 9. The features are sorted by mean rank. Grayed-out features were unavailable for the given task. The first four tasks are evaluated on ranking predictions on CodeT5 transfer experiments, while code generation is evaluated over ranking predictions on Llama 3.3 70B-Instruct.
@@
While Section 4.1 explored how well transfer learning performs for different languages on different tasks, we would like to dig deeper into why they perform that way. That is, how do the features of language pairs affect transfer? Which characteristics of the source and target language best explain the variation of transfer across tasks, the dependence of transfer on source and target pairing, as well as the superior transfer supported by specific sources? These questions can be answered by measuring feature importance.
@@
The state-of-the-art technique for measuring feature importance, SHAP values (Lundberg & Lee, 2017), considers all possible subsets $S\subseteq F$ of features. Then, the importance of a given feature is the difference it makes to model performance in combination with all possible sets of other features. In the context of ranking different source languages for cross-lingual transfer, SHAP values measure how different features contribute to transfer. Figure 8 shows SHAP values computed based on the ranking models evaluated in Section 4.2. For presentation purposes, it rescales the values so the most important feature for each task has importance 1.
Task-dependent feature importance.
@@
While previous works on transfer learning across programming languages emphasize the importance of specific features for cross-lingual transfer, the lack of comparison on different tasks limits their conclusions. The supported cross-task analysis demonstrates a task-dependent feature importance whereby different features contribute differently to cross-lingual transfer depending on the task. For example, the top features for the solution domain classification task are Difficulty, Literals, Names, and Operators. Being able to predict the solution domain for a code solution requires some knowledge of the underlying problem for which the tags are an attribute. The difficulty score is another attribute of the problem. Problems with similar difficulty scores require similar algorithms. The literals, names, and operators are other indicators of the algorithms used in a code sample. In comparison, the most significant features for the clone detection task are Token, Names, and Keywords. Detecting a clone requires different skills than classifying domains. A deeper understanding of the code semantics irrespective of the similar algorithms used is needed. The overlaps in names, keywords, and more generally tokens is key for understanding the semantics of code (Ahmed & Devanbu, 2022).
Prior work on transfer learning in natural languages (Lin et al., 2019) showed feature importances to vary greatly across tasks. We confirm this for programming languages.
Range of important features.
@@
Different tasks seem to not only focus on different features, but also on a different number of features. While the feature importance for solution domain classification and error detection is concentrated on a small number of features, it is more spread out for clone detection and code repair. For example, the code repair task requires a transfer of knowledge from overlapping keywords, names, and more generally different tokens from a source language for fixing a code in a target language. On the other hand, solution domain classification focuses on fewer features, related to the problems attributes.
Higher-level tasks seem to draw upon more features to best understand the code.
Comparison across categories.
@@
Comparing the importance of different feature categories across tasks, we observe a relatively higher significance of syntactic features as demonstrated by the top two features Keywords and Names. Compared to transfer learning across natural languages, model-specific and linguistics features seem to carry less significance for transfer in programming languages. While de Vries et al. (2022) find the potential of cross-lingual transfer dropping significantly for sources and targets that are not seen by the model, our finding brings hope to not only transferring learning to low-resource target languages, but even finetuning the model on unseen languages that could serve as good sources, such as Kotlin.
Overall, the most important features across the four tasks are keywords and names.
Within-category differences.
@@
While different categories seem to have a higher impact on the model predictions, the impact of single features within a category varies across task. Our results are consistent with prior work that found Names and Keywords to be the most important features for cross-lingual transfer in certain tasks (Ahmed & Devanbu, 2022). However, half of the tasks show keywords to be more important than names, while the other half has the opposite order of importance.
4.4 Few-shot prompting
We examine cross-lingual transfer learning for programming languages by pairing low- and high-resource languages across classification and code-repair tasks. Whereas our initial study employed CodeT5-base, we now replicate a representative subset of experiments with the larger Llama 3.3 70B-Instruct model to test the generalizability of our conclusions to contemporary large-scale LLMs. For evaluation we adopt the few-shot prompting scheme described in Section 3; this technique has been shown to boost out-of-domain code-generation quality for previously unseen languages and to markedly reduce the compilation and parsing failures that arise when the model lacks explicit exposure to a target language (Athiwaratkun et al., 2023).
We investigate whether rankers constructed from the features introduced earlier can accurately predict the best source language(s) for few-shot prompting. We compare different rankers explored in the previous section and test whether our ranker model can generalize to larger models. We train a ranker model on the same features described in Table 9 on Pass@1 scores. Figure 7 shows mean NDCG@3 scores. The learned ranker still surpasses other rankers and heuristic baselines, underscoring that considering multiple features is essential for accurately identifying the top source languages. SHAP analysis further reveals that expression of types and overlap in literals dominate the other features in predicting top performing sources. Languages whose type systems are richly expressed in code provide stronger semantic signals that transfer directly for code generation of other similarly typed languages solutions. Meanwhile a high lexical overlap in literals such as numeric and string constants offers immediate token-level anchors that reduce out-of-vocabulary risk during generation. Together, these two features jointly encode semantic structure and surface-level similarity, making them the most dependable indicators of best source languages for code generation.
5 Limitations
This section describes the limitations of our study, including the rationale for key design choices and the steps we took to address those limitations.
5.1 External validity
One limitation is that, to keep the large number of experiments feasible and reduce their carbon footprint, we constrained our fine-tuning experiments to a moderately sized model. We address this by using CodeT5 (Wang et al., 2021), a commonly adopted model in AI-for-Code research. Even if a larger model were used, insights from one model would not necessarily generalize to another. Accordingly, we included model-specific features in our study to account for potential model dependence. These features are described in Table 9 and analyzed as part of Section 4.3. Despite model-specific features, we do acknowledge that experimenting with different models, not just of varying size but also architecture, would have made our study more robust. Because the source and target languages are usually different and come in many combinations, we do not necessarily run the risk of overfitting. However, using models of varying strength could have exposed a broader range of more interesting patterns which would have made our study richer and more general. To mitigate this limitation, we conduct separate experiments with a larger model and few-shot prompting.
Another limitation is that, to cover multiple languages, we limited ourselves to only five tasks. We address this by selecting tasks that span a broad range of difficulty and demonstrate diversity, as indicated by their differing performance and feature importance. Nonetheless, our study includes more tasks than any previous cross-lingual transfer study on programming languages.
@@ 5.1 External validity
A further limitation is that our datasets include synthetic elements—for instance, type-IV clones (Roy & Cordy, 2007) in the clone detection data or fault injection in the code repair data. However, using exclusively natural data imposes a demanding constraint that few academic studies meet, and it is also uncommon in industrial practice. To address this, we explicitly measure the importance of dataset-specific features for each language pair.
@@
Another limitation is that we focused on the setting of zero data in the target language. This represents the lowest barrier to entry, as gathering even a small number of examples can be tricky. Therefore, we defer exploring scenarios that include some target-language data to future work.
5.2 Internal validity
@@
We believe that our study avoids one common internal validity limitation, which revolves around insufficient runs. Across all experiments, each source-target pair is evaluated only once. But thanks to the sheer number of language pairs for each task, we get thousands of different experiments, whose results are consistent with each other. Moreover, each of these experiments involves fine-tuning with tens of thousands of samples, giving us confidence that the choice of samples is not a limitation either. However, having results based on single runs does limit our ability to asses variability in experiments. Since our study is the first of its kind at this scale, it lacks directly comparable baselines. We mitigate this limitation by noting that, wherever comparisons are possible, our results corroborate those of narrower studies—for instance, prior work on transfer for identifier-centric tasks (Ahmed & Devanbu, 2022).
6 Conclusion
We perform a systematic and extensive study of LLM transfer learning covering up to 41 programming languages, across 5 tasks including both classification and generation. Programming languages covered include several low-resource but often still widely-used languages. Cross-lingual transfer learning performs much better than zero-shot with an LLM that has been pre-trained on code. One interesting finding is that even a language not seen during pre-training, like Kotlin, can be good fine-tuning source language across several target languages and tasks. On the other hand, certain languages that are often used extensively to pre-train LLMs, like C++ and Python, are worse source or target languages relative to others.
To understand relative differences in cross-lingual performance between different programming languages, we define several linguistic, syntactic, dataset, and model-specific features of language pairs and then analyze the feature importance for a model that predicts transfer performance. We show how different features are needed for ranking source languages, compared to ad-hoc heuristics such as overlap on names. To explain how a language that was unseen during pre-training, like Kotlin, can be a good source language, we show that seen language features are less significant compared to dataset, linguistic, and syntactic source language features. Similar to previous work of Ahmed & Devanbu (2022), we show that keywords and names are top features on average for CodeT5. On the other hand, unlike previous work, we cover more languages and more diverse tasks, and find that feature importance vary strongly across tasks. Replicating a representative subset of experiments with a larger model under few-shot prompting confirms that these feature hierarchies and the performance prediction gains they enable generalize to state-of-the-art LLMs. Overall, we believe that this paper sheds light on how learning transfers among programming languages, and that this ultimately leads to better models to assist users of those languages, particularly low-resource ones.
7 Data Availability
The experiments are based on the publicly available CodeT5-base (220M parameters) model (Wang et al., 2021) and the open-sourced datasets CodeNet (Puri et al., 2021) and XCodeEval (Khan et al., 2023). Our code is available at https://github.com/baltaci-r/XPL.